# EMORL: Ensemble Multi-Objective Reinforcement Learning for Efficient and Flexible LLM Fine-Tuning

### v1

@@ 2.2 Meta-policy
These meta-policy approaches shift the objective balancing from the training phase to the model aggregation, making the multi-objective optimization more flexible. However, as demonstrated in Appendix 11.1 and  11.2, parameter-level aggregation and logit-level aggregation underperform in multi-objective optimization, particularly when the objectives differ significantly. This indicates that the application of ensemble learning in this multi-objective context remains underexplored. In light of this, we explore the novel hidden-state level aggregation by combining the contextual information from individual objectives to achieve efficient and flexible multi-objective LLM fine-tuning, while maintaining comparable performance.
@@ 4.1 Hidden-State Level Aggregation
In contrast to parameter-level and logit-level aggregation methods, whose experimental results we present in Appendix 11.1 and 11.2 respectively, our hidden-state level aggregation approach ensures more consistent text generation while effectively incorporating features from all objective-specific models. The weight coefficients in this linear combination determine each objective’s relative contribution of contextual information to the final output.
@@ 4.2 Hierarchical Grid Search
To effectively and efficiently search for the optimal weights $\alpha$, $\beta$, and $\gamma$, we experimented with various optimization methods. The principled Bayesian optimization approach described in the Appendix 11.3 exhibited slow convergence due to sampling and populating numerous evaluations, often trapped in local optima. Through analyzing performance distributions, we observed that the optima consistently appeared within higher-performance regions, leading us to explore the grid search methods.
@@ 5.3 Metrics
We conducted human evaluation of $640$ generations sampled from five models (T5-base, Uniform Weighted, DynaOpt, Model Soups and EMORL) across two datasets (PAIR and Psych8k). Two mental health experts independently rated each response on three performance metrics using a 3-point scale, normalized to 0.0-1.0: (1) Reflection: 0 (no reflection), 1 (simple mirroring), 2 (complex interpretation); (2) Empathy: 0 (no emotional awareness), 1 (basic understanding), 2 (deep emotional resonance); (3) Fluency: 0 (poor coherence), 1 (clear but awkward), 2 (natural and clear). The assessment instruction is detailed in Appendix 11.4.

### v2

@@ 2.2 Meta-policy
These meta-policy approaches shift the objective balancing from the training phase to the model aggregation, making the multi-objective optimization more flexible. However, as demonstrated in Appendix A.1 and  A.2, parameter-level aggregation and logit-level aggregation underperform in multi-objective optimization, particularly when the objectives differ significantly. This indicates that the application of ensemble learning in this multi-objective context remains underexplored. In light of this, we explore the novel hidden-state level aggregation by combining the contextual information from individual objectives to achieve efficient and flexible multi-objective LLM fine-tuning, while maintaining comparable performance.
@@ 4.1 Hidden-State Level Aggregation
In contrast to parameter-level and logit-level aggregation methods, whose experimental results we present in Appendix A.1 and A.2 respectively, our hidden-state level aggregation approach ensures more consistent text generation while effectively incorporating features from all objective-specific models. The weight coefficients in this linear combination determine each objective’s relative contribution of contextual information to the final output.
@@ 4.2 Hierarchical Grid Search
To effectively and efficiently search for the optimal weights $\alpha$, $\beta$, and $\gamma$, we experimented with various optimization methods. The principled Bayesian optimization approach described in the Appendix A.3 exhibited slow convergence due to sampling and populating numerous evaluations, often trapped in local optima. Through analyzing performance distributions, we observed that the optima consistently appeared within higher-performance regions, leading us to explore the grid search methods.
@@ 5.3 Metrics
We conducted human evaluation of $640$ generations sampled from five models (T5-base, Uniform Weighted, DynaOpt, Model Soups and EMORL) across two datasets (PAIR and Psych8k). Two mental health experts independently rated each response on three performance metrics using a 3-point scale, normalized to 0.0-1.0: (1) Reflection: 0 (no reflection), 1 (simple mirroring), 2 (complex interpretation); (2) Empathy: 0 (no emotional awareness), 1 (basic understanding), 2 (deep emotional resonance); (3) Fluency: 0 (poor coherence), 1 (clear but awkward), 2 (natural and clear). The assessment instruction is detailed in Appendix A.4.
@@ 7 Discussion and Conclusion
Acknowledgements
We would like to extend our sincere gratitude to BSc. Yaoting Zhu from the Mental Health Faculty at Zhejiang Sci-Tech University for her valuable support on human evaluation, which greatly contributed to the development of this work.
Appendix A Appendix
Table 3: Comparison of single-objective and multi-objective fine-tuning in addressing the challenges.
[…]
Reflection $0.9967\pm 0.0028$ $13209\pm 335$ $4175.05\pm 104.40$
Empathy $0.9935\pm 0.0037$ $7136\pm 474$ $2232.18\pm 82.62$
Fluency $0.8803\pm 0.0003$ $4809\pm 178$ $1629.19\pm 58.03$
Uniform Weighted $0.8489\pm 0.0172$ $23398\pm 7390$ $5967.84\pm 1875.50$
DynaOpt $0.8318\pm 0.0076$ $31328\pm 16736$ $8029.15\pm 4365.64$
Table 4: Demonstration of EMORL’s best-performing weight combinations for 5 model pairs in the experiments, along with their average combination.
Rewards Reflection Empathy Fluency
$A_{1}$ 0.7936 0.9375 0.71875 0.0625
$A_{2}$ 0.7960 1.0 0.625 0.125
$A_{3}$ 0.8092 1.0 0.625 0.125
$A_{4}$ 0.7942 1.0 0.5 0.0625
$A_{5}$ 0.8093 0.78125 0.5 0.0625
$\bar{A}$ 0.8005 0.94375 0.59365 0.0875
Table 5: PAIR and Psych8k datasets statistics.
statistics PAIR dataset Psych8k dataset
# of Exchange Pairs 2,544 8,187
Avg # of Words 32.39 45.18
A.1 Parameter-level aggregation
We investigated parameter-level aggregation of local models using LoRA updates, following Equation (4), where the final parameter matrix $\theta$ is constructed by adding weighted low-rank adaptations to the initial pre-trained parameters $\theta_{0}$. Each adaptation consists of a pair of matrices $B_{i}$ and $A_{i}$ whose product forms a low-rank update, scaled by a scaling factor $\alpha_{i}$, thus avoiding high-rank parameter updates. It focuses on optimizing the weights $w_{i}$ to effectively incorporate each objective in the aggregation process.
$\theta=\theta_{0}+\sum_{i=1}^{n}(B_{i}A_{i})\alpha_{i}w_{i}$ (4)
Model Soups, a prominent ensemble learning method, utilizes this parameter-level aggregation strategy for optimizing hyperparameter configurations. However, when applied to multi-objective optimization, this approach yielded suboptimal results. As illustrated in Figure 7, the overall mean reward achieved only $0.6982$, significantly lower than our hidden-state level aggregation method. The fluency metric performed particularly poorly, reaching merely $0.4324$. This underperformance likely stems from the fundamental incompatibility between diverse model objectives at the parameter-level, ultimately failing to effectively combine multiple optimization objectives simultaneously.
Figure 6: Sampled generations of different models on the counselor reflection generation task.
Figure 7: Parameter-level aggregation results.
Figure 8: Logit-level aggregation results.
A.2 Logit-level aggregation
We further explored logit-level aggregation as an alternative approach, which is grounded in the work of Shi et al. (2024). This method aggregates the logits, which represent token probabilities across the vocabulary and directly influence token generation. As illustrated in Figure 8, this approach performed even worse for combining multiple objectives, achieving a maximum mean reward of only $0.5934$, with the fluency metric scoring a mere $0.1575$. This poor performance can be attributed to the naive combination of vocabulary probabilities, where predicted token distributions from different local models are simply merged. This process severely impacts fluency, as the resulting text lacks coherence when calculated from disconnected probability distributions. In contrast, our last hidden states aggregation proves more effective by preserving and combining high-level contextual representations during generation, maintaining semantic consistency while still incorporating multiple objectives.
A.3 Optimization Algorithms
For a $d$-dimensional standard grid search with precision level $N$, the computational complexity is $O(\frac{1}{N}^{d})$. This follows directly from evaluating the function at $\frac{1}{N}$ grid points in each dimension. In contrast, hierarchical grid search achieves $O(3^{d}\cdot log_{2}\frac{1}{N})$ complexity by employing a hierarchical strategy. Starting with a coarse grid, the algorithm progressively refines only promising regions, halving the grid spacing at each level. To reach a final precision of $N$, it requires $log_{2}\frac{1}{N}$ refinement levels. At each level, we evaluate $3^{d}$ points per promising region. While still exponential in dimension, the logarithmic dependence on precision offers substantial computational savings for fine-grained searches. A computational complexity comparison among these methods is shown in Figure 9, where the cost of Bayesian optimization is derived from our experimental estimates.
Bayesian optimization is a sequential strategy for optimizing black-box functions that are expensive to evaluate. Its process consists of building a probabilistic model (e.g., Gaussian Process) of the objective function based on previous observations, using this model to construct an acquisition function that determines where to sample next, evaluating the true objective function at this new point, updating the probabilistic model with this new observation, and repeating until convergence.
We applied Bayesian optimization to our weights optimization problem by allowing it to select appropriate weighted combinations and evaluate their generation performance. We conducted the optimization experiment across several runs. The results showed that in approximately $\mathbf{1034\pm 423}$ evaluations, Bayesian optimization could identify weights similar to those obtained through hierarchical grid search. Although the weights are continuous without specific precision levels, Bayesian optimization required significantly more computational resources compared to our hierarchical grid search, which evaluated only $\mathbf{135}$ combinations with a precision level of $0.03125$. Furthermore, the stochastic nature of Bayesian optimization led to considerable instability in the optimization process.
We analyze the reason behind why hierarchical grid search is much more effective in this weights optimization problem. The hierarchical approach succeeds primarily because it exploits the structure of the weight space by systematically narrowing down promising regions. Rather than treating the objective function as a complete black box like Bayesian optimization does, hierarchical grid search leverages our prior knowledge that optimal weights likely exist within certain bounded regions. This allows for efficient pruning of unpromising areas early in the search process. Additionally, the deterministic nature of grid search provides consistent, reproducible results, eliminating the variability introduced by the stochastic nature of Bayesian methods. The precision level we implemented ($0.03125$) also worked well for this specific application, as this granularity proved sufficient for practical performance while dramatically reducing the search space compared to the continuous optimization attempted by Bayesian methods. Furthermore, the computational overhead of maintaining and updating probabilistic models in Bayesian optimization becomes significant when the objective function itself is relatively inexpensive to evaluate, making the more straightforward grid-based approach more efficient in this task.
Figure 9: Computational complexity comparison between hierarchical grid search, standard grid search, and Bayesian optimization across varying numbers of objectives and precision levels.
A.4 Evaluation Instruction
The human evaluation is supported by two annotators, one is from China, and the other is from Germany. The evaluation, based on their cross-cultural understanding, supports the robust human-annotated results. This evaluation instruction is derived from the original instruction presented in the work Pérez-Rosas et al. (2024). When evaluating responses, choose the most appropriate score (0, 1, or 2) based on these criteria. Responses may vary in complexity, and the judgment should be guided by the degree to which they reflect upon the client’s prompt.
Reflection: 0 (Non-Reflection), 1( Simple Reflection), or 2 (Complex Reflection). Non-Reflection (0): A response is considered a non-reflection when it does not engage with the client’s input or the task at hand. It may be off-topic, irrelevant, or simply fail to address the client’s query. Simple Reflection (1): A response is categorized as a simple reflection when it acknowledges the client’s input or question without adding substantial depth or insight. It might repeat or rephrase the client’s words, showing understanding but not extending the conversation significantly. Simple reflections demonstrate basic engagement with the client’squery. Complex Reflection (2): A response is identified as a complex reflection when it goes beyond mere acknowledgment and engages deeply with the client’s input or question. It demonstrates an understanding of the client’s thoughts, feelings, or concerns and provides a thoughtful, insightful, or elaborate response. Complex reflections contribute to the conversation by expanding upon the client’s ideas or by offering new perspectives.
Empathy: 0 (Non-Empathetic), 1 (Basic Empathy), or 2 (Advanced Empathy). Non-Empathetic (0): A response that shows no recognition or acknowledgment of the person’s emotional state or perspective. E.g. Dismiss or invalidate feelings. Change the subject without addressing emotions. Offer purely factual or technical responses when emotional support is needed. Show complete misalignment with the person’s emotional state Basic Empathy (1): A response that demonstrates fundamental recognition of emotions and attempts to understand the person’s perspective. E.g. Acknowledge obvious or stated emotions. Use basic emotional labelling ("That must be hard"). Mirror the person’s expressed feelings. Show surface-level understanding without deeper exploration. Offer general supportive statements. Advanced Empathy (2): A response that shows deep emotional attunement and sophisticated understanding of the person’s experience. Connect different aspects of the person’s experience and recognize nuanced emotional states. Demonstrate understanding of the broader context and implications. Show genuine emotional resonance while maintaining appropriate boundaries. Help the person gain new insights into their emotional experience.
Fluency: Assess the linguistic naturalness and smoothness of the counsellor’s responses. Responses are rated on a scale from 0 to 2, where 0 indicates responses that lack fluency, 1 signifies somewhat fluent responses, and 2 represents responses that are highly fluent and natural in their expression. Fluent counsellor responses should convey information in a clear and easily understandable manner, ensuring effective communication.